# Convergence and error control of consistent PINNs for elliptic PDEs

### v1

@@ 8 A numerical optimal recovery algorithm
Let ${\cal U}$ be the model class (3.3) determined by the model classes ${\cal F}$ and ${\cal G}$, where ${\cal F}$ is a maximal model class (3.24), (3.26) and ${\cal G}$ is a maximal model class (3.28). To build a numerical algorithm for the recovery of $u\in{\cal U}_{\rm data}$, we want to minimize the discrete loss ${\cal L}^{*}$ over a suitable set $\Sigma_{n}$. Let $M\geq 1$ and let $\widetilde{u}$ be such that
@@
$\|u-\widetilde{u}\|_{H^{1}(\Omega)}\lesssim M{\cal R}_{{\cal U}}(\tilde{m},\bar{m}),\quad m\geq 1.$ (8.2)
with the implied constant depending only on $r$ and $d$. In other words, $\widetilde{u}$ provides a near optimal recovery when $d\geq 3$ and a near optimal recovery up to a $\log m$ factor when $d=2$. Of course, the minimization problem (8.1) is taken over too large of a set to be numerically viable and so we would like to utilize minimization over a smaller set.
Let $\Sigma_{n}\subset H^{1}(\Omega)$ be an approximating set. Here, the primary examples for $\Sigma_{n}$ are linear spaces of finite dimension $n$ or a nonlinear manifolds depending on $n$ parameters such as NNs. For $M\geq 1$ (typically taken as some fixed number not dependent on $n$) we define the set
@@
and consider the solution $\widehat{u}$ of the minimization problem
@@
We assume for the moment that $d\geq 3$ and show that $\widehat{S}$ is a near optimal recovery of $u$ from the given data provided the set $\Sigma_{n}(M)$ is sufficiently good at approximating the elements of our model class ${\cal U}$. Later, we explain the changes needed for the case $d=2$.
@@
$E(v,\Sigma_{n}(M)):=\begin{cases}\inf_{S\in\Sigma_{n}(M)}\left(\|\Delta v-\Delta S\|_{L_{\gamma}(\Omega)}+\|Tr(v-S)\|_{H^{1/2}(\partial\Omega)}\right)&d\geq 3,\\ \\ \inf_{S\in\Sigma_{n}(M)}\left((1+\log(\widetilde{m}))\|\Delta v-\Delta S\|_{L_{\gamma}(\Omega)}+\|Tr(v-S)\|_{H^{1/2}(\partial\Omega)}\right)&d=2.\end{cases}$ (8.5)
the error of simultaneously approximating $v$ and $\Delta v$. We remark that in the case $d=2$ and $1<p\leq\infty$, then this analysis can be modified to remove the logarithm. Then, for the model class ${\cal U}$ of $H^{1}(\Omega)$, we denote by
@@
the error of restricted approximation of the class ${\cal U}$ by elements from $\Sigma_{n}(M)$.
We have the following theorem.
@@
with the constant in $\lesssim$ depending only on $r,d$. Moreover, if the approximation set $\Sigma_{n}(M)$ is such that $E({\cal U},\Sigma_{n}(M))\leq CR_{{\cal U}}(\widetilde{m},\overline{m})$ for some $C$ depending only on $r,d$, then $\widehat{S}$ provides a uniform near optimal recovery (up to a logarithmic factor in the case $d=2$) for the model class ${\cal U}$.
@@
${\cal L}^{*}(\widehat{S})\lesssim E(u,\Sigma_{n}(M))+M{\cal R}_{{\cal U}}(\widetilde{m},\overline{m}).$ (8.8)
When this is inserted in (8.7) we obtain the theorem. ∎
9 CPINNs
The previous section has shown that any minimizer of the loss ${\cal L}^{*}$ over $\Sigma_{n}(M)$ provides a near optimal recovery of $u$ provided $\Sigma_{n}(M)$ is sufficiently good at approximating the elements of the model class ${\cal U}$. In this section, we specialize these results to the case where $\Sigma:=\Sigma_{n}$ is a space generated by neural networks with $n$ parameters and a prescribed architecture. As a primary example, we can take the activation function $\sigma$ to be ${\rm ReLU}^{k}$ with $k\geq 3$. This guarantees that the elements in $\Sigma_{n}$ have continuous Laplacians, i.e., $\Delta S$ is continuous whenever $S\in\Sigma_{n}$. This also guarantees that ${\cal L}^{*}(S)$ is well defined for each $S\in\Sigma_{n}$. Regarding the architecture of the network, we consider the case of deep networks which is the space usually utilized in PINNs. A similar discussion applies to shallow networks.
Consistent PINNs (CPINNs): We call any algorithm which minimizes the loss ${\cal L}^{*}$ over a neural network space $\Sigma_{n}$ consistent PINNs and denote it by the acronym CPINNs. The discussion in Remark 7.6 has given the advantages of using CPINNs over standard PINNs with the loss ${\cal L}$. Here, we do not necessarily require the minimization take place over the constrained space $\Sigma_{n}(M)$, unless explicitly stated since the latter invokes numerical difficulties in implementation.
We do not address the question of designing an optimization algorithm that is guaranteed to come close to minimizing the loss ${\cal L}^{*}$ over $\Sigma_{n}$ or $\Sigma_{n}(M)$. We know that the lack of such theoretical guarantees is a bottleneck in rigorously proving convergence results. We analyze the performance of any minimizer or near minimizer of the loss without addressing how such a minimizer is found.
@@ 9 CPINNs
Let $\widehat{u}_{n}$ denote any minimizer of (8.4). Theorem 8.1 guarantees that $\widehat{u}_{n}$ is a near optimal recovery of $u$ provided that for the model class ${\cal U}$, we can make $E({\cal U},\Sigma_{n}(M))$ small by taking $n$ large. Thus, the question is whether $\Sigma_{n}$ has favorable restricted approximation properties. There is a large literature of approximation theoretic results concerning how efficiently deep neural networks can approximate various classes of functions (see [19] and the references therein). Let us briefly discuss the specific case of approximating functions from Besov spaces using deep ReLU neural networks. When error is measured in the $L_{p}$-norm, optimal rates have been obtained [55, 64, 52]. There has also been recent work on the problem when error is measured in another Besov space, although in this case a smoother activation function must be taken [63, 62].
On the other hand, the question of restricted approximation that is used in the present paper has seemingly not been studied. We conjecture that the techniques developed in [55, 64, 52] can be appropriately modified to give the same rates for restricted approximation with a fixed suitably large value for $M$ (independent of $n$). Since the proofs are already quite difficult, we leave the proof of this conjecture to future work.
10 Numerical Illustrations
In this section, we present some numerical experiments in support of the theoretical results of the previous sections. The main take away from these theoretical results is that when using optimization of a prescribed loss function over a set $\Sigma_{n}$ to numerically compute an approximation $\widehat{u}$ to the solution $u$ of (1), then the loss function ${\cal L}^{*}$ given by (1.9) as in CPINNs is the proper choice. The purpose of the numerical experiments of this section is to analyze the effect of using ${\cal L}^{*}$ rather than the standard loss ${\cal L}$ when taking $\Sigma_{n}$ as a neural network space.
As a prelude to presenting our numerical experiments, we make some remarks on minimization of loss functions over $\Sigma_{n}$ that are well known to practitioners. Since both losses are non convex when viewed as a function of the parameters of the neural network space $\Sigma_{n}$, there is not a known mathematical procedure which is proven to converge to a minimizer of the loss. Typical approaches used in practice, try various initial guesses of the parameters and use variations of gradient descent with variable step sizes, i.e., learning rates, in an attempt to find a minimizer of the loss. There is no a priori guarantee that any specific recipe will work. In other words, one cannot blindly apply any prescribed procedure with an a priori guarantee of success.
Another drawback of this approach to solving the PDE is that one cannot test its efficacy unless the true solution is in some sense known so that one can measure performance a posteriori. One advantage of using the loss ${\cal L}^{*}$ in place of ${\cal L}$ is that our theory provides an a posteriori bound of the error (see (7.3)) that does not require knowledge of the solution $u$. In other words, when using ${\cal L}^{*}$, one can guarantee an a posterior bound of the error by computing ${\cal L}^{*}(S)$, where $S$ is the output of the numerical optimization procedure. Indeed, one would compute $\|S\|_{\mathcal{U}}$ and from (7.3) we know
$\|u-S\|_{H^{1}(\Omega)}\lesssim\mathcal{L}^{*}(S)+(1+\|S\|_{{\cal U}})R_{{\cal U}}(\widetilde{m},\overline{m}).$ (10.1)
Here the second term is the recovery error. All quantities on the right side of (10.1) only involve $S$ and so one has a bound for the error without knowing $u$.
Let us also note that in practice, for the PDE setting and in other learning scenarios, one does not try to impose the model class restriction in the optimization routine since it is computationally too expensive. We follow this paradigm in our numerical procedures.
Finally, we want to point out that the minimization of the new loss $\mathcal{L}^{*}$ differs from the standard loss $\mathcal{L}$ in advocating the use of:
  • a discrete $L_{\gamma}(\Omega)$ norm (we take $\gamma=1.1$ when $d=2$) rather than a discrete $L_{2}(\Omega)$ for the PDE residual;
  • a discrete $H^{1/2}(\partial\Omega)$ norm rather than a discrete $L_{2}(\partial\Omega)$ norm for the boundary condition residual.
Although it is not possible to completely separate the effect of these two modifications, we consider two different test cases corresponding to a harmonic solution $u$ with non-vanishing boundary conditions and a singular solution with vanishing boundary conditions; see Sections 10.2. In each case, we compare the results of the minimization of the new loss function (referred as CPINNs) and the standard loss (referred as PINNs).
10.1 Practical Algorithm
We now describe the practical algorithm implemented for the minimization of the losses ${\cal L}^{*}$ and ${\cal L}$. We remark that PINNs algorithms rely on the inherent assumption that the minimization processes are successful. As noted above, there is no prescribed recipe for the initial choice of parameters, the activation functions, the optimization algorithm and the hyperparameters that will guarantee convergence. Rather, one is left to experimenting with different choices and employing different heuristics until satisfactory convergence of this method is obtained.
There is no way to explore all of the possible minimization algorithms. Moreover, once a particular numerical optimization procedure is chosen to implement, one is always subjected to the possibility that some other choice would have been better and demonstrated different results. So,our numerical results must be viewed from this perspective.
Our approach in this section is to find one procedure that works reasonably well in minimizing the loss (1.9). We remark that in order to conform the common practice of minimizing the squared residual in PINNs, we actually used a loss function obtained by squaring each of the terms in (1.9) and (1.8) in our experiments, i.e. we minimize the losses ($sq$ is shorthand for squared)
${\cal L}_{sq}^{*}:=\!\!\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}|f({\bf x}_{i})+\Delta v({\bf x}_{i})|^{\gamma}\right]^{2/\gamma}\!\!\!\!+\left[\frac{1}{\overline{m}^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\overline{m}}\frac{|[g-v]({\bf z}_{i})-[g-v]({\bf z}_{j})|^{2}}{|{\bf z}_{i}-{\bf z}_{j}|^{2}}\right]\!\!\!+\left[\frac{1}{\overline{m}}\sum_{j=1}^{\overline{m}}|g({\bf z}_{j})-v({\bf z}_{j})|^{2}\right],$
and
${\cal L}_{sq}:=\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}[\Delta v({\bf x}_{i})+f({\bf x}_{i})]^{2}\right]+\left[\frac{1}{\overline{m}}\sum_{i=1}^{\overline{m}}[v({\bf z}_{i})-g({\bf z}_{i})]^{2}\right].$
@@ 10.1 Practical Algorithm
Here we take a uniform grid of collocation points and $\widetilde{m}$ is the total number of collocation points (where we penalize the residual of the equation), while $\overline{m}$ is the number of collocation points which lie on the boundary (where we penalize the residual of the boundary values).
We then apply the exact same procedure to minimize both the original PINNs loss ${\cal L}_{sq}$ and the CPINNs loss ${\cal L}_{sq}^{*}$. We observe from the final loss values that our procedure obtains sufficient minimization of both loss functions, so that our results enable us to compare the effect of using the two different loss functions ${\cal L}_{sq}$ and ${\cal L}_{sq}^{*}$. These results demonstrate a significant improvement in accuracy when minimizing the loss ${\cal L}_{sq}^{*}$ instead of ${\cal L}_{sq}$.
Finally, we remark that the running time of the training the we have implemented is very high and is not remotely competitive with traditional numerical methods for this problem. Improving the efficiency of the training of PINNs on both ${\cal L}_{sq}$ and ${\cal L}_{sq}^{*}$ is an important and active area of research (see for instance [44, 14, 37, 48, 66]), and incorporating these recently developed techniques into our algorithm is an area of future work.
@@
In our experiments, we use a deep neural network with the following structure. We use a deep network with $L=8$ hidden layers of width $W=100$. The first layer uses a $\tanh$ activation function, while the remaining layers are residual layers [33] with the ReLU3 activation function. The network is initialized as follows:
@@
  • In the final layer the weights are initialized randomly from a normal distribution with variance $1/\sqrt{W}$, which ensures that the network output has the same variance as the output of the final hidden layer.
We train the networks with gradient descent with an initial step size of $10^{-3}$ decreasing by a factor of $2$ every $4000$ steps and momentum $0.9$ for a total of $40,000$ steps. We also rescale the velocities (i.e. weighted averages of the previous gradients according to the momentum) in each layer to have the same maximum as the parameters in that layer, we have found that this makes the training somewhat more stable and improves the final loss values obtained. The full code and all experimental details can be found at https://github.com/jwsiegel2510/consistent-PINNs.
We also note that although the $H^{1/2}(\partial\Omega)$ norm is nonlocal, the implementation of the discrete $H^{1/2}(\partial\Omega)$-norm present in ${\cal L}_{sq}^{*}$ can easily be vectorized and is fairly efficient in our practical implementation. Of course, the computation and optimization of the discrete $H^{1/2}(\partial\Omega)$-norm scales quadratically with the number of collocation points on the boundary, while the norm used in the standard loss ${\cal L}_{sq}$ scales only linearly. In our examples, the number of collocation points in the interior scales quadratically with the number of points on the boundary, so that using the discrete $H^{1/2}(\partial\Omega)$-norm on the boundary is comparably expensive to the penalty on the interior. In addition, the evaluation of the neural network at the collocation points typically dominates the calculation, rather than the calculation of the norms or their gradients from these values. However, when using a large number of collocation points on the boundary optimizing the new loss function ${\cal L}_{sq}^{*}$ may be less efficient than optimizing the original ${\cal L}_{sq}$ loss. This can potentially be mitigated by using a truncated version of the discrete $H^{1/2}(\partial\Omega)$-norm, but we leave this to future work.
@@ 10.2 Results
In our first experiment, we solve the Poisson equation
$-\Delta u=f~{}\text{in $\Omega$},~{}u=g~{}\text{on $\partial\Omega$},$ (10.2)
where $f$ and $g$ are chosen so that the exact solution is given by
$u(x,y)=e^{x}\cos(\pi y),\qquad(x,y)\in\Omega=(0,1)^{2}.$
For this solution $u$, we have $f=0$ and the boundary values $g$ are non-zero. The results of this experiment with different numbers of collocation points $\widetilde{m}$ can be found in Table 1.
In our second experiment, we solve the Poisson equation (10.2) where $f$ and $g$ are chosen so that the exact solution is given by

### v2

@@ 9 A numerical optimal recovery algorithm
Let ${\cal U}$ be the model class (3.3), determined by ${\cal F}$ and ${\cal G}$, where ${\cal F}$ is a maximal model class (3.27), (3.29) and ${\cal G}$ is a maximal model class (3.31). To build a numerical algorithm for the recovery of $u\in{\cal U}_{\rm data}$, we want to minimize the discrete loss ${\cal L}^{*}$ over a suitable set $\Sigma_{n}$. Let $M\geq 1$ and let $\widetilde{u}$ be such that
@@
$\|u-\widetilde{u}\|_{H^{1}(\Omega)}\lesssim M{\cal R}_{{\cal U}}(\tilde{m},\bar{m}),\quad m\geq 1,$ (9.2)
with the implied constant depending only on $r$ and $d$. In other words, $\widetilde{u}$ provides a near optimal recovery when $d\geq 3$ and a near optimal recovery up to a $\log m$ factor when $d=2$. However, the minimization problem (9.1) is taken over too large of a set to be numerically viable and so we would like to utilize minimization over a smaller set.
Let $\Sigma_{n}\subset H^{1}(\Omega)$ be an approximating set. Here, the primary examples for $\Sigma_{n}$ are linear spaces of finite dimension $n$ or a nonlinear manifolds depending on $n$ parameters such as NNs. For $M\geq 1$ (typically taken as some fixed number not dependent on $n$), we define the set
@@
and consider the solution $\widehat{S}$ to the minimization problem
@@
$E(v,\Sigma_{n}(M)):=\begin{cases}\inf_{S\in\Sigma_{n}(M)}\left(\|\Delta v-\Delta S\|_{L_{\gamma}(\Omega)}+\|Tr(v-S)\|_{H^{1/2}(\partial\Omega)}\right),&d\geq 3,\\ \\ \inf_{S\in\Sigma_{n}(M)}\left((1+\log(\widetilde{m}))\|\Delta v-\Delta S\|_{L_{\gamma}(\Omega)}+\|Tr(v-S)\|_{H^{1/2}(\partial\Omega)}\right),&d=2,\end{cases}$ (9.5)
the error of simultaneously approximating $v$ and $\Delta v$. We remark that in the case $d=2$ and $1<p\leq\infty$, this analysis can be modified to remove the logarithm. Then, for the model class ${\cal U}$ of $H^{1}(\Omega)$, we denote by
@@
the error of restricted approximation of the class ${\cal U}$ by elements from $\Sigma_{n}(M)$. We have the following theorem.
@@
with the constant in $\lesssim$ depending only on $r,d$. Moreover, if the approximation set $\Sigma_{n}(M)$ is such that $E({\cal U},\Sigma_{n}(M))\leq C{\cal R}_{{\cal U}}(\widetilde{m},\overline{m})$ for some $C$ depending only on $r,d$, then $\widehat{S}$ provides a uniform near optimal recovery (up to logarithmic factors in the case $d=2$) for the model class ${\cal U}$.
@@
${\cal L}^{*}(\widehat{S})\lesssim E(u,\Sigma_{n}(M))+M{\cal R}_{{\cal U}}(\widetilde{m},\overline{m}),$ (9.8)
which together with (9.7) proves the theorem. ∎
Remark 9.2.
@@
Theorem 9.1 guarantees that $\widehat{S}$ is a near optimal recovery of $u$ provided that for the model class ${\cal U}$ we can make $E({\cal U},\Sigma_{n}(M))$ small by taking $n$ large. Note that the quantity $E({\cal U},\Sigma_{n}(M))$ determines how well the elements of the set $\Sigma_{n}(M)$ simultaneously approximate the Laplacians (in the $L_{\gamma}(\Omega)$ norm) and the traces (in $H^{1/2}(\partial\Omega)$ norm) of the elements in ${\cal U}$. Thus, the question is whether $\Sigma_{n}$ has favorable restricted approximation properties. There is a large literature of approximation theoretic results concerning how efficiently deep neural networks can approximate various classes of functions (see [24] and the references therein). When error is measured in the $L_{p}$-norm, optimal rates have been obtained [65, 77, 62] in the case of approximating functions from Besov spaces using deep ReLU neural networks. There has also been recent work on the problem when error is measured in another Besov space, although in this case a smoother activation function must be taken [76, 75]. On the other hand, the question of restricted approximation that is used in the present paper has seemingly not been studied. We conjecture that the techniques developed in [65, 77, 62] can be appropriately modified to give the same rates for restricted approximation with a fixed suitably large value for $M$ (independent of $n$).
10 Numerical Experiments
In this section, we present numerical experiments testing the loss functions described in §8 for the practical solution of elliptic PDEs. In our experiments, we solve the two-dimensional Poisson equation
$-\Delta u=f~{}\text{in $\Omega$},~{}u=g~{}\text{on $\partial\Omega$},$ (10.1)
@@ 10 Numerical Experiments
on $\Omega=(0,1)^{2}$ for various right-hand sides $f$ and boundary values $g$. We take a uniform grid of collocation points on $\Omega$, where $\widetilde{m}$ is the total number of collocation points (where we penalize the residual of the equation), and $\overline{m}$ is the number of collocation points which lie on the boundary (where we penalize the residual of the boundary values). We then test the following four loss functions, described in §8:
  • The original (unweighted) PINNs loss: ${\cal L}_{sq,1}$;
  • A properly weighted PINNs loss: ${\cal L}_{sq,\lambda(\overline{m})}$, see (8.5);
  • The loss ${\cal L}^{*}_{sq,\gamma}$ from (7.2) (since $d=2$ we choose here $\gamma=1.1$);
  • The loss ${\cal L}^{*}_{sq,2}$ (the loss function (7.2) with $L_{2}$ on the domain).
Since we are in the case $d=2$, the latter two loss functions involve a logarithm which is ignored in the numerical experiments. For all losses, we use the same neural network and training algorithm. We report both the final loss values achieved and the final relative error in the $H^{1}(\Omega)$-norm (estimated using a large number of collocation points).
Let us remark on the relative cost of implementing each of these four loss functions. In all our experiments, the efficiency of optimizing each of them was roughly the same. The reason for this is that although the discretization of the $H^{1/2}$-norm on the boundary used in the losses ${\cal L}^{*}_{sq,\gamma}$ and ${\cal L}^{*}_{sq,2}$ is more expensive to compute and differentiate, the training process is dominated by the cost of evaluating the network and its derivatives at all of the collocation points rather than evaluating the loss based upon these values. For this reason, using losses which discretize the $H^{1/2}$-norm doesn’t significantly slow down the network training.
10.1 Network Architecture and Training Algorithm
When training PINNs, it is known that the optimization of the neural network can be particularly sensitive and tricky [60, 46, 71, 72, 18], and much research has been devoted to the development of improved training algorithms for PINNs [53, 17, 44, 57, 79, 70]. In this section, we describe our network architecture, initialization, and training algorithm in detail.
@@ 10.1 Network Architecture and Training Algorithm
In our experiments, we use a deep neural network with $L$ hidden layers of width $W$. The first layer uses a $\tanh$ activation function, while the remaining layers are residual layers [40] with the ReLU3 activation function. The network is initialized as follows:
@@
  • In the final layer, the weights are initialized randomly from a normal distribution with variance $1/\sqrt{W}$, which ensures that the network output has the same variance as the output of the final hidden layer.
We train this network applying the energy natural gradient descent algorithm described in [57], using a line search to find the step size in each iteration. The full code and all experimental details can be found at https://github.com/jwsiegel2510/consistent-PINNs.
@@ 10.2 Results
Experiment 1: In our first experiment, we choose $f$ and $g$ so that the exact solution is given by
$u(x,y)=e^{x}\cos(y),\qquad(x,y)\in\Omega=(0,1)^{2}.$
The function is harmonic, i.e $f\equiv 0$. The function $u$ is very smooth and so for this problem it suffices to use a relatively small number of collocation points and a relatively small network. We test a grid of collocation points with $5,10,15$ and $20$ points in each direction. We use a small network ($L=3$, $W=5$) and iterate for $500$ steps. We present the results in Table 1.
$\widetilde{m}$ $\overline{m}$ ${\cal L}_{sq,1}$ ${\cal L}_{sq,1}$ ${\cal L}_{sq,\lambda(\overline{m})}$ ${\cal L}_{sq,\lambda(\overline{m})}$ ${\cal L}^{*}_{sq,\gamma}$ ${\cal L}^{*}_{sq,\gamma}$ ${\cal L}^{*}_{sq,2}$ ${\cal L}^{*}_{sq,2}$
Rel Error Loss Rel error Loss Rel error Loss Rel error Loss
25 16 $0.80\%$ $2.8\cdot 10^{-5}$ $0.68\%$ $1.1\cdot 10^{-4}$ $0.33\%$ $2.5\cdot 10^{-5}$ $0.30\%$ $3.1\cdot 10^{-5}$
100 36 $0.43\%$ $1.7\cdot 10^{-5}$ $0.21\%$ $2.3\cdot 10^{-5}$ $0.21\%$ $2.0\cdot 10^{-5}$ $0.25\%$ $2.2\cdot 10^{-5}$
225 56 $0.41\%$ $1.5\cdot 10^{-5}$ $0.15\%$ $2.1\cdot 10^{-5}$ $0.23\%$ $1.7\cdot 10^{-5}$ $0.27\%$ $2.0\cdot 10^{-5}$
400 76 $0.41\%$ $1.7\cdot 10^{-5}$ $0.086\%$ $7.18\cdot 10^{-5}$ $0.22\%$ $1.9\cdot 10^{-5}$ $0.33\%$ $3.6\cdot 10^{-5}$
Table 1: Experimental results on the harmonic problem. All relative errors are in the $H^{1}(\Omega)$-norm, computed using a grid with $500$ points in each direction. When the number of grid points is large, the properly weighted PINNs loss ${\cal L}_{sq,\lambda(\overline{m})}$ results in significantly lower error than the other loss functions.
Experiment 2: In our second experiment, we choose $f$ and $g$ so that the exact solution is given by
$u(x,y)=\frac{e^{2(x+y)}\cos(2\pi(y-x))}{1+8x^{2}+y^{2}},\qquad(x,y)\in\Omega=(0,1)^{2}.$
In this case, both $f$ and $g$ are non-zero and the solution is significantly less smooth. To capture this, we use a slightly larger network ($L=3$ and $W=10$). All other parameters are the same as in the first experiment. The results are presented in Table 2.
$\widetilde{m}$ $\overline{m}$ ${\cal L}_{sq,1}$ ${\cal L}_{sq,1}$ ${\cal L}_{sq,\lambda(\overline{m})}$ ${\cal L}_{sq,\lambda(\overline{m})}$ ${\cal L}^{*}_{sq,\gamma}$ ${\cal L}^{*}_{sq,\gamma}$ ${\cal L}^{*}_{sq,2}$ ${\cal L}^{*}_{sq,2}$
Rel Error Loss Rel error Loss Rel error Loss Rel error Loss
25 16 $26\%$ $1.2\cdot 10^{-7}$ $24\%$ $2.2\cdot 10^{-6}$ $24\%$ $3.3\cdot 10^{-7}$ $22\%$ $1.6\cdot 10^{-7}$
100 36 $1.5\%$ $3.7\cdot 10^{-3}$ $2.9\%$ $3.7\cdot 10^{0}$ $1.1\%$ $1.6\cdot 10^{-2}$ $0.69\%$ $2.0\cdot 10^{-2}$
225 56 $1.4\%$ $5.5\cdot 10^{-3}$ $1.1\%$ $4.1\cdot 10^{-2}$ $1.6\%$ $1.1\cdot 10^{-1}$ $0.90\%$ $1.1\cdot 10^{-2}$
400 76 $2.3\%$ $1.7\cdot 10^{-2}$ $0.88\%$ $3.2\cdot 10^{-2}$ $1.5\%$ $4.7\cdot 10^{-2}$ $0.75\%$ $2.2\cdot 10^{-2}$
Table 2: Experimental results on the smooth problem. All relative errors are in the $H^{1}(\Omega)$-norm, computed using a grid with $500$ points in each direction. Aside from large errors when the number of collocation points is relatively small, the consistent loss with $L_{2}$-penalty on the domain, ${\cal L}^{*}_{sq,2}$, performs best on this experiment.
Experiment 3: Finally, for our last experiment, we choose $f$ and $g$ so that the exact solution is given by